OpenStudy (anonymous):

What is the formula for the volume of a cone with radius r and a height of 2r?

OpenStudy (anonymous):

V=πr2h 3

OpenStudy (johnweldon1993):

Normal Volume:\[\large \pi r^2 \frac{h}{3}\] if 'r' = r and h = 2r we have \[\large \pi r^2\frac{2r}{3}\] \[\large \frac{2\pi r^3}{3}\]

OpenStudy (anonymous):

https://www.google.com/#q=volume+of+a+cone Google can calculate volume of cones